# DanModCap: Designing a Danmaku Moderation Tool for Video-Sharing Platforms that Leverages Impact Captions

### v3

@@ Abstract.
Content Moderation, User Engagement, Danmaku, Large Language Models, Emotional Resonance, Cognitive Resonance, Pro-social Behavior
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Collaborative and social computing systems and toolsccs: Human-centered computing Collaborative and social computing design and evaluation methods
Figure 1. Our Danmaku moderation approach. 1) Typical Danmaku involves positive, inappropriate, or irrelevant content. 2) On video-sharing platforms, Danmaku is a built-in social feature that supports end-users’ social communication and is synchronized with the video. 3) To moderate Danmaku content, we introduce a proactive approach called Impact Captions, that leverages a key feature in East Asian variety shows. 4) Building on this conceptual foundation, we introduce DanModCap, a tool for Danmaku content moderation that uses generative AI to create Impact Captions. 5 and 6) The pipeline of Impact Caption generation and example captions, which include the Tsukkomi style (red), the Explanatory style (blue), and the Humorous and Praising style (orange).
@@ 1. Introduction
  • The design and implementation of DanModCap, a Danmaku content moderation system for VSP that integrates Impact Captions.
  • A nuanced understanding of the role of cognitive and emotional resonance in VSP content moderation.
@@ 3.1.2. Caption Style 2: Multi-Perspective Information
With Multi-Perspective Information, two perspectives were used: first person point of view (POV) and third person POV. First person POVs centered on the protagonist’s perspective, directly disclosing their experiences, thoughts, actions, and emotions. These captions were instrumental in forging an empathetic connection between the lead character and the audience. In contrast, third person POVs offered a more comprehensive, ‘removed’ viewpoint, encompassing additional details, commentaries, and humorous response styles like Tsukkomi (Roasting), Expository or Humorous and Praise. This approach significantly broadened the entertainment value and deepened audience understanding, offering an enriched viewing experience.
@@ 5.1. Design Goals
  • DG1: Support the proactive moderation of Danmaku through textual and non-textual Impact Caption elements to improve the efficiency of curation and reduce the need for post-video processing.
  • DG2: Encourage spontaneous prosocial user behavior for Danmaku-streaming content to reduce moderation burdens and improve the overall user experience.
  • DG3: Construct cognitive and emotional resonance via textual and non-textual elements while being mindful of relevance, cultural sensitivity, and visual appeal.
@@ 5.2. Tenants of Impact Captions
UTF8gbsn The Tsukkomi (Roasts) style is utilized with negatively oriented Danmaku content. It leverages an informal communication approach, commonly identifying loopholes or keywords in another’s language or behavior, followed by the expression of irony or doubt. According to interpersonal communication theory, this style employs humor to address critical comments while avoiding excessive negativity, thus preventing the treatment of viewers as “emotional garbage” (Hofmann et al., 2016). For instance, hyperbolic metaphors might be used to humorously address narcissistic tendencies, like “这位自恋狂是不是以为自己是全宇宙的中心啊?你有皇位要继承吗? (Does this narcissist think they are the center of the universe? Do they have a throne to inherit?)”. This style facilitates the collective criticism of disruptive Danmaku content, promoting emotional regulation through shared complaints, which can engender a sense of unity and intimacy among strangers (Suls et al., 2002; Ellemers et al., 2002).
The Expository style is targeted at Danmaku that are irrelevant or ambiguous and serves to redirect the viewer’s attention back to the video. The use of metaphors related to ’adventures’ or ’deciphering,’ such as ”前方高能!(High energy ahead.)” or ”路飞的尼卡太阳神形态来了!(Here comes Luffy in his Sun God Nika form!)”, assists in clarifying the video content for the viewer.
Lastly, the Humorous and Praise style echoes and amplifies positive Danmaku content to foster a pleasant atmosphere and evoke positive emotions. It is useful in contexts that warrant enjoyable commentary, such as compliments. For example, in a vibrant travel vlog, employing phrases like ”神仙UP主 (God-tier UP)” or ”优雅!实在是太优雅了!(Elegant! Truly, utterly elegant!)” can create a lively Danmaku environment. This style enhances viewer engagement and strengthens the emotional connection between the viewer and the content, ultimately leading to positive viewer responses (John and Gross, 2004).
@@ 6. DanModCap Evaluation
  • RQ3: How does DanModCap impact users’ perception of the Danmaku, the community, and VSPs?

### v4

@@ Abstract.
Content Moderation, Danmaku, Video Sharing Platform, Emotional Resonance, Cognitive Resonance, Human-AI Co-Creation, Pro-social Behavior, Large Language Models
copyright: acmlicensedjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 2article: CSCW056publicationmonth: 4doi: 10.1145/3710954ccs: Human-centered computing Collaborative and social computing systems and toolsccs: Human-centered computing Collaborative and social computing design and evaluation methods
Figure 1. Our Danmaku moderation approach. 1) Typical Danmaku involves positive, inappropriate, or irrelevant content. 2) On video-sharing platforms, Danmaku is a built-in social feature that supports end-users’ social communication and is synchronized with the video. 3) To moderate Danmaku content, we introduce a proactive approach called Impact Captions, that leverages a key feature in East Asian variety shows. 4) Building on this conceptual foundation, we introduce DanModCap, a tool for Danmaku content moderation that uses generative AI to create Impact Captions. 5 and 6) The generation pipeline and example captions.
@@ 1. Introduction
  • The design and implementation of DanModCap, a Danmaku content moderation system for video-sharing platforms that integrates Impact Captions.
  • A nuanced understanding of the role of cognitive and emotional resonance in video-sharing platforms content moderation.
@@ 3. Study 1: Impact Caption Video Analysis
3.1. Procedure
We initially consulted three global internet film and television databases, i.e., IMDB [^1] IMDB: https://www.imdb.com/, TV Time [^2] TV Time: https://www.tvtime.com/, and Douban [^3] Douban: https://www.douban.com/, to identify shighly popular variety shows from East Asia that had an airing history of at least eight years and had consistently incorporated Impact Caption content. These included China’s Who is the Murderer [^4] Who is the Murderer: https://en.wikipedia.org/wiki/Who%27s_the_Murderer, South Korea’s Running Man [^5] Running Man: https://en.wikipedia.org/wiki/Running_Man_(TV_video), and Japan’s VS ARASHI [^6] VS ARASHI: https://zh.wikipedia.org/wiki/VS%E5%B5%90. We then watched the top three episodes from each show that had the highest viewership to gain a preliminary understanding of the show’s content. While watching, we identified that Impact Captions had several diverse visual design features that influenced audience appeal and manipulation based on previously work (Sasamoto, 2014). For example, we found that videos with Impact Caption employed a series of striking visual design elements, including vivid color contrasts, dynamic typography layouts, and theme-coordinated background images.
Following this, we examined YouTube [^7] YouTube: https://www.youtube.com/ and Bilibili [^8] Bilibili: https://www.bilibili.com/ to analyze current design trends and the utilization of Impact Captions in variety shows on these platforms as online video platforms allow for more experimentation and innovation in visual design compared to more constrained television formats. Three researchers conducted keyword searches on these video-sharing platforms (e.g., “variety show Impact Caption”, “Impact Caption in the TV program” and “video Impact Caption”) to find and select representative samples based on the number of views. After filtering out irrelevant and duplicate content, 27 videos with an average duration of 43.23 minutes and 33.83k views were selected for analysis. The researchers then conducted independent open coding of a subset of the videos to identify critical aspects of Impact Captions, including visual elements, interaction patterns, and the perspective, using methods established in prior work ( (Sasamoto, 2014; Lu and Lu, 2022)). Following multiple rounds of reflection and discussion, a taxonomy of Impact Captions was developed, with three types of captions that spanned seven design dimensions (Figure 2). The outcomes from this analysis were used to inform our subsequent research activities.
@@ 3.2.2. Caption Style 2: Multi-Perspective Information
With Multi-Perspective Information, two perspectives were used: first person points of view (POV) and third person POV. First person POVs centered on the protagonist’s perspective, directly disclosing their experiences, thoughts, actions, and emotions. These captions were instrumental in forging an empathetic connection between the lead character and the audience. In contrast, third person POVs offered a more comprehensive, ‘removed’ viewpoint, encompassing additional details, commentaries, and humorous response styles like Tsukkomi (Roasting), Expository or Humorous and Praise. This approach significantly broadened the entertainment value and deepened audience understanding, offering an enriched viewing experience.
@@ 5.1. Design Goals
  • Design Goal 1: Support the proactive moderation of Danmaku through textual and non-textual Impact Caption elements to improve the efficiency of curation and reduce the need for post-video processing.
  • Design Goal 2: Encourage spontaneous prosocial user behavior for Danmaku-streaming content to reduce moderation burdens and improve the overall user experience.
  • Design Goal 3: Construct cognitive and emotional resonance via textual and non-textual elements while being mindful of relevance, cultural sensitivity, and visual appeal.
@@ 5.2. Tenants of Impact Captions
UTF8gbsn
5.2.1. Generation Style 1:
The Tsukkomi (Roasts) style is utilized with negatively oriented Danmaku content. It leverages an informal communication approach, commonly identifying loopholes or keywords in another’s language or behavior, followed by the expression of irony or doubt. According to interpersonal communication theory, this style employs humor to address critical comments while avoiding excessive negativity, thus preventing the treatment of viewers as “emotional garbage” (Hofmann et al., 2016). For instance, hyperbolic metaphors might be used to humorously address narcissistic tendencies, like “这位自恋狂是不是以为自己是全宇宙的中心啊?你有皇位要继承吗? (Does this narcissist think they are the center of the universe? Do they have a throne to inherit?)”. This style facilitates the collective criticism of disruptive Danmaku content, promoting emotional regulation through shared complaints, which can engender a sense of unity and intimacy among strangers (Suls et al., 2002; Ellemers et al., 2002).
5.2.2. Generation Style 2:
The Expository style is targeted at Danmaku that are irrelevant or ambiguous and serves to redirect the viewer’s attention back to the video. The use of metaphors related to ‘adventures’ or ‘deciphering,’ such as “前方高能!(High energy ahead.)” or “路飞的尼卡太阳神形态来了!(Here comes Luffy in his Sun God Nika form!)”, assists in clarifying the video content for the viewer.
5.2.3. Generation Style 3:
The Humorous and Praise style echoes and amplifies positive Danmaku content to foster a pleasant atmosphere and evoke positive emotions. It is useful in contexts that warrant enjoyable commentary, such as compliments. For example, in a vibrant travel vlog, employing phrases like “神仙UP主 (God-tier UP)” or “优雅!实在是太优雅了!(Elegant! Truly, utterly elegant!)” can create a lively Danmaku environment. This style enhances viewer engagement and strengthens the emotional connection between the viewer and the content, ultimately leading to positive viewer responses (John and Gross, 2004).
@@ 6. DanModCap Evaluation
  • RQ3: How does DanModCap impact users’ perception of the Danmaku, the community, and Video-Sharing Platforms?
@@ 9. Conclusion
10. Acknowledgement
{CJK*}
UTF8gbsn We are thankful to all reviewers for their constructive feedback, which significantly strengthened this paper. We also gratefully acknowledge the inspiration drawn from various variety shows (particularly “Who is the Murderer 明星大侦探”) and video-sharing platforms, which sparked new ideas and perspectives for this research. In addition, we extend our appreciation to our lab-mates for their insightful discussions and support throughout the research and review process.